# Block-to-Scene Pre-training for Point Cloud Hybrid-Domain Masked Autoencoders

### v1

@@ 4 Experiments
First, we pre-train the PointHDMAE model using our block-to-scene pretraining strategy based on point cloud data from the scene domain. After pre-training, we directly transfer the pre-trained model to various downstream tasks in different point cloud domains for fine-tuning. During fine-tuning, we selectively activate different sub-modules depending on the domain of the point cloud; for instance, we activate the object encoder for object point clouds, while utilizing all encoders for scene point clouds. This strategy enables our PointHDMAE model, pre-trained with the block-to-scene approach, to outperform existing domain-specific models in most cases without any additional domain adaptation training, demonstrating the generalization capability of our model.
4.1 Pre-training
@@
Datasets. We combined all training data from the two most commonly used indoor scene datasets, SUNRGB-D (Song et al., 2015) and ScanNetV2 (Dai et al., 2017), to construct our pretraining dataset. Specifically, SUNRGB-D includes 5K single-view RGB-D training samples with oriented bounding box annotations for 37 object categories. ScanNetV2 contains 1.2K training samples, each with axis-aligned bounding box labels belonging to 18 object categories. We extracted 50K points for each of the 6.2K samples, using only the xyz coordinates of each point to construct the pretraining dataset.
Table 1: Classification accuracy on real-scanned (ScanObjectNN) and synthetic (ModelNet40) point clouds. In ScanObjectNN, we report the overall accuracy (%) on three variants. In ModelNet40, we report the overall accuracy (%) for both without and with voting. "#Params" represents the model’s parameter count.
Method Reference #Params (M) ScanObjectNN ModelNet40
OBJ-BG OBJ-ONLY PB-T50-RS w/o Vote w/ Vote
Supervised Learning Only
PointNe (Qi et al., 2017a) CVPR 2017 3.5 73.3 79.2 68 89.2 -
PointNet++ (Qi et al., 2017b) NeurIPS 2017 1.7 82.3 84.3 77.9 90.7 -
DGCNN (Wang et al., 2019) TOG 2019 1.8 82.8 86.2 78.1 92.9 -
PointMLP (Ma et al., 2022) ICLR 2022 12.6 - - 85.2 94.1 94.5
SFR (Zha et al., 2023a) ICASSP 2023 - - - 87.8 93.9 -
P2P-HorNet (Wang et al., 2022) NeurIPS 2022 195.8 - - 89.3 94.0 -
Single Modal Self-Supervised Learning
Point-BERT (Yu et al., 2022) CVPR 2022 22.1 87.43 88.12 83.07 92.7 93.2
MaskPoint (Liu et al., 2022) ECCV 2022 22.1 89.30 88.10 84.30 - 93.8
Point-MAE (Pang et al., 2022) ECCV 2022 22.1 90.02 88.29 85.18 93.2 93.8
Point-M2AE (Zhang et al., 2022) NeurIPS 2022 15.3 91.22 88.81 86.43 93.4 94.0
PointGPT-S (Chen et al., 2024) NeurIPS 2023 29.2 91.63 90.02 86.88 - 94.0
PointDif (Zheng et al., 2024) CVPR 2024 - 93.29 91.91 87.61 - -
MaskFeat3D (Yan et al., 2024) ICLR 2024 15.3 93.20 91.50 88.40 - 94.0
PointMamba (Liang et al., 2024) NeurIPS 2024 12.3 94.32 92.60 89.31 93.6 -
PointHDMAE (Ours) Ours 22.8 95.18 93.12 90.01 93.7 94.2
Multimodal Self-Supervised Learning
Joint-MAE (Guo et al., 2023) IJCAI 2023 - 90.94 88.86 86.07 - 94.0
ACT (Dong et al., 2023) ICLR 2023 22.1 93.29 91.91 88.21 93.2 93.7
TAP+PointMLP (Wang et al., 2023) ICCV 2023 12.6 - - 88.50 94.0 -
I2P-MAE (Zhang et al., 2023) CVPR 2023 15.3 94.15 91.57 90.11 93.7 94.1
Recon (Qi et al., 2023) ICML 2023 44.3 95.18 93.29 90.63 94.1 94.5
@@
Pre-training. During the pretraining phase, we input all 50K×3 point clouds into the scene encoder of PointHDMAE to extract scene-level features. Simultaneously, we randomly select 32 point blocks from the scene point cloud, with each block containing 2K local points, and input these into the 32 object encoders with shared parameters. We use the AdamW (Kingma & Ba, 2014) optimizer with a base learning rate of 5e-4 and a weight decay of 0.1. Simple rotation is applied as data augmentation to both the scene point cloud and each selected object point cloud. For the object point cloud masks, we set the mask ratio to 60% following previous work (Pang et al., 2022; Dong et al., 2023). We train the model from scratch for 200 epochs using 8 A100 GPUs. Furthermore, we can also leverage pre-trained object point cloud models on the ShapeNet55 (Chang et al., 2015) dataset to initialize our object-level models.
@@ 4.2.1 Object Point Cloud Classification
We first evaluate the performance of our model on object point cloud classification tasks using the object encoder of PointHDMAE. We conduct point cloud classification on two of the most commonly used object point cloud datasets: ScanObjectNN (Uy et al., 2019) and ModelNet40 (Wu et al., 2015). ScanObjectNN contains 15K real scanned point clouds, each with various backgrounds, occlusions, and noise, which effectively assesses the model’s robustness. ModelNet40 includes 12K synthetic point clouds belonging to 40 different categories, with each point cloud being complete and clean, providing a better representation of 3D object shapes.
Following previous work (Dong et al., 2023; Liang et al., 2024), we use 2K points as input for ScanObjectNN, apply simple rotation for data augmentation, and report classification accuracy without using voting. For ModelNet40, we use 1K points as input, apply scale and translate data augmentation, and report classification accuracy both without voting and with the standard voting mechanism.
As presented in Table 1, firstly, compared to the recent state-of-the-art method PointMamba, our approach surpasses it by 0.86%, 0.52%, and 0.70% on the three variants of ScanObjectNN, respectively. This improvement is significant, given that the same downstream task settings were used. Secondly, our method surpasses the majority of multimodal pre-trained models, ranking just the same with the leading Recon (Qi et al., 2023). This is still highly competitive as Recon (Qi et al., 2023) benefits from the supplementary knowledge of image, and language modalities, while also requiring significantly more parameters than our method. Our PointHDMAE achieves leading performance indicating that using randomly selected point blocks from the scene, despite lacking explicit real-world significance, can still be effectively used for mask reconstruction pre-training. This effectiveness arises because mask reconstruction primarily focuses on learning general representations through the reconstruction of the original shapes, without requiring a specific understanding of the shapes’ concrete meanings.
@@ 4.2.2 Scene Point Cloud Detection
We further fine-tune the pre-trained PointHDMAE on scene-level object detection tasks. At this stage, we primarily rely on the scene encoder to process scene-level inputs. Simultaneously, we randomly select 32 point blocks from the scene point cloud and use the 32 shared object encoders to handle these local point blocks. During the decoding phase, we integrate the results from the scene encoder with random queries from the scene, helping the scene encoder to better understand scene details. We use ScanNetV2 (Dai et al., 2017), to evaluate our model’s scene understanding capabilities.
Table 4.2.2 shows our detection results, our PointHDMAE model has shown significant improvements compared with other models. For example, compared to the previous state-of-the-art domain-specific pretraining model PointDif (Zheng et al., 2024), our method achieves a 6.2% improvement on $AP_{50}$. This substantial improvement is mainly attributed to our PointHDMAE using multiple object encoders to assist the scene encoder in analyzing the overall scene. This approach enables the scene model to focus on more local details, thereby enhancing scene understanding.
@@
Methods $AP_{25}$ $AP_{50}$ Supervised Learning Only VoteNet (Qi et al., 2019) 58.6 33.5 3DETR (Misra et al., 2021) 62.1 37.9 Single Modal Self-Supervised Learning PointContrast (Xie et al., 2020b) 58.5 38.0 STRL (Huang et al., 2021) 59.5 38.4 DepthContrast (Zhang et al., 2021) 61.3 - Point-BERT (Yu et al., 2022) 61.0 38.3 MaskPoint (Liu et al., 2022) 64.2 42.1 PointDif (Zheng et al., 2024) - 43.7 PointHDMAE (Ours) 66.8 49.9 Multimodal Self-Supervised Learning PiMAE (Chen et al., 2023) 62.6 39.4 ACT (Dong et al., 2023) 63.8 42.1 TAP (Wang et al., 2023) 62.6 39.4
Methods $\mathrm{mIoU}_{c}$ $\mathrm{mIoU}_{I}$ Supervised Learning Only PointNet++ (Qi et al., 2017b) 81.9 85.1 PointMLP (Ma et al., 2022) 84.6 86.1 Single-Modal Self-Supervised Learning PointContrast (Xie et al., 2020b) - 85.1 CrossPoint (Afham et al., 2022) - 85.5 Point-BERT (Yu et al., 2022) 84.1 85.6 MaskPoint (Liu et al., 2022) 84.4 86.0 Point-MAE (Pang et al., 2022) 84.2 86.1 PointGPT-S (Chen et al., 2024) 84.1 86.2 PointMamba (Liang et al., 2024) 84.4 86.2 PointHDMAE (Ours) 85.0 86.3 Multimodal Self-Supervised Learning ACT (Dong et al., 2023) 84.7 86.1 Joint-MAE (Guo et al., 2023) 85.4 86.3 Recon (Qi et al., 2023) 84.8 86.4
4.2.3 Object Point Cloud Part Segmentation
We assess the performance of our PointHDMAE in part segmentation using the ShapeNetPart dataset (Chang et al., 2015), comprising 16,881 samples across 16 categories. We utilize the same segmentation setting after the pre-trained encoder as in previous works Pang et al. (2022); Zhang et al. (2022) for fair comparison. The experimental results, displayed in Table 4.2.2, demonstrate that our model exhibits competitive performance in tasks such as part segmentation, which demands a more fine-grained understanding of point clouds.
4.2.4 Object Point Cloud Completion
Previous pretraining models have not explored the effects on low-level tasks. We first specifically design task heads for downstream low-level point cloud completion tasks. Then, we fine-tune our PointHDMAE model on point cloud completion on the classic point cloud completion dataset PCN (Yuan et al., 2018) and ShapeNet-55 (Chang et al., 2015). The PCN dataset is created from the ShapeNet (Chang et al., 2015) dataset, including eight categories with a total of 30974 CAD models. Compared to the PCN dataset, ShapeNet-55 includes 55 different categories of 3D models. Following previous practices (Yu et al., 2021; Li et al., 2023), we used 41,952 models for training and 10,518 models for testing. For each object, we randomly sampled 8,192 points from the surface to obtain the point cloud. We also divide the test samples into three difficulty degrees, simple, moderate, and hard in our experiments and we report the performance for each method in simple, moderate, and hard to show the ability of each network to deal with tasks at difficulty levels.
We followed the data processing methods established in previous works (Yu et al., 2021; Li et al., 2023) and report the average $l_{1}$ Chamfer Distance (Fan et al., 2017) (CD-Avg) across all 8 object categories in the PCN dataset, and the average Chamfer Distance for the 55 categories under the simple (CD-S), moderate (CD-M), and hard (CD-H) settings in the ShapeNet-55 dataset, along with the overall average of these three metrics (CD-Avg). As shown in table 4, our model achieves state-of-the-art results across various settings in both datasets, demonstrating that our pre-trained model can better handle diverse scenarios, such as different viewpoints, object categories, incomplete patterns, and varying levels of incompleteness, even in lower-level completion tasks.
Table 4: Quantitative comparison of point cloud completion task on PCN. Point resolutions for the output and ground-truth are 16384. For Chamfer Distance, lower is better.
Methods Reference PCN ShapeNet-55
CD-Avg CD-S CD-M CD-H CD-Avg
FoldingNet (Yang et al., 2018) CVPR 2018 14.31 2.67 2.66 4.05 3.12
PCN (Yuan et al., 2018) 3DV 2018 9.64 1.94 1.96 4.08 2.66
GRNet (Xie et al., 2020a) ECCV 2020 8.83 1.35 1.71 2.85 1.97
PoinTr (Yu et al., 2021) ICCV 2021 8.38 0.58 0.88 1.79 1.09
LAKeNet (Tang et al., 2022) CVPR 2022 7.23 - - - 0.89
SnowFlakeNet (Xiang et al., 2021) ICCV 2021 7.21 0.70 1.06 1.96 1.24
ProxyFormer (Li et al., 2023) CVPR 2023 6.77 0.49 0.75 1.55 0.93
SeedFormer (Zhou et al., 2022) ECCV 2022 6.74 0.50 0.77 1.49 0.92
PointHDMAE Ours 6.54 0.51 0.70 1.24 0.81
4.2.5 Scene Semantic Segmentation
@@
We have conducted experiments on scene-level semantic segmentation tasks to assess the performance of PointHDMAE in classifying each point in a scene into semantic categories. We validated our model using the indoor S3DIS (Armeni et al., 2016) dataset for semantic segmentation tasks. Specifically, we tested the model on Area 5 while training on other areas and report the mean IoU (mIoU) and mean Accuracy (mAcc). To ensure a fair comparison, we used the same codebase based on the PointNext (Qian et al., 2022a) baseline and employed identical decoders and semantic segmentation heads. We acknowledge that there are several outstanding works in semantic segmentation, such as PointTransformerV3 (Wu et al., 2024), that achieve performance far exceeding that of PointNeXt. However, these advanced models often require more complex inputs, such as color and normal information. Our focus is on segmentation using only the most basic xyz coordinate information. Therefore, we chose PointNeXt as the baseline codebase for our pretraining model and ensure a fair comparison with other pretraining models.
@@
The experimental results are shown in the table 4.2.5. Compared to training the PointNeXt model from scratch, our method improves the mIoU score by 2.3%. It also shows significant improvements over other pretraining models, such as Point-MAE (Pang et al., 2022) and PointDif (Zheng et al., 2024). This enhancement is largely due to our block-to-scene pretraining, which equips the model with strong scene understanding capabilities and further demonstrates the generalizability of our approach.
@@
Methods mIoU mAcc PointCNN (Li et al., 2018) 57.3 63.9 Pix4Point (Qian et al., 2022b) 69.6 75.2 PointNeXt (Qian et al., 2022a) 68.5 75.1 Point-BERT (Yu et al., 2022) 68.9 76.1 MaskPoint (Liu et al., 2022) 68.6 74.2 Point-MAE (Pang et al., 2022) 68.4 76.2 PointDif (Zheng et al., 2024) 70.0 77.1 PointHDMAE (Ours) 70.8 77.6
Table 6: Head tuning of object-level classification and scene-level detection.
Object-Level Classification Methods OBJ-BG OBJ-ONLY Point-BERT (Yu et al., 2022) 88.81 88.3 Point-MAE (Pang et al., 2022) 89.50 88.98 PointHDMAE (Ours) 90.71 88.47 Scene-Level Detection Methods $AP_{25}$ $AP_{50}$ Point-BERT (Yu et al., 2022) 60.13 38.5 Point-MAE (Pang et al., 2022) 60.82 40.4 PointHDMAE (Ours) 62.44 41.8
4.2.6 Head Tuning on Downstream Tasks
We further demonstrate the generalization capability of our pre-trained model by performing only head tuning on downstream tasks. We keep the pre-trained feature extractor fixed and only train the task-specific heads (including the classification head and detection head). At the same time, all models use the same downstream task setting as described in 4.2. These experiments are designed to better evaluate the performance of the proposed PointHDMAE and to facilitate a more direct comparison with the baselines. The results of these experiments are summarized in the tables 4.2.5. As shown, our proposed Point-HDMAE still demonstrates superior performance compared to the baselines, highlighting its robust representational capabilities even without fine-tuning the entire model.

### v3

@@ 4 Experiments
First, we pre-train the Point-MoDE model using our block-to-scene pretraining strategy based on point cloud data from the scene domain. After pre-training, we directly transfer the pre-trained model to various downstream tasks in different point cloud domains for fine-tuning. During fine-tuning, we selectively activate different sub-modules depending on the domain of the point cloud; for instance, we activate the object expert encoder for object point clouds, while utilizing all encoders for scene point clouds. This strategy enables Point-MoDE model, pre-trained with the block-to-scene approach, to outperform existing domain-specific models in most cases without any additional domain adaptation training, demonstrating the superiority and transferability of our model.
4.1 Dataset and Pre-training
@@
We initialize our object expert model using the Point-MAE or other models pre-trained on ShapeNet Chang et al. (2015), which can be directly obtained from these official repositories. Then, We combined all training data from the two most commonly used indoor scene datasets, SUNRGB-D Song et al. (2015) and ScanNetV2 Dai et al. (2017), to construct our block-to-scene pretraining dataset. Specifically, SUNRGB-D includes 5K single-view RGB-D training samples with oriented bounding box annotations for 37 object categories. ScanNetV2 contains 1.2K training samples, each with axis-aligned bounding box labels belonging to 18 object categories. We extracted 50K points for each of the 6.2K samples, using only the xyz coordinates of each point to construct the pretraining dataset.
@@
During the pretraining phase, we input all 50K×3 point clouds into the scene encoder of Point-MoDE to extract scene-level features. Simultaneously, we randomly select 32 point blocks from the scene point cloud, with each block containing 2K local points, and input these into the 32 object encoders with shared parameters. We use the AdamW optimizer with a base learning rate of 5e-4 and a weight decay of 0.1. Simple rotation is applied as data augmentation to both the scene point cloud and each selected object point cloud. For the object point cloud masks, we set the mask ratio to 60% following previous work Pang et al. (2022); Dong et al. (2023). We train the model for 200 epochs using 8 A100 GPUs.
@@ 4.2.1 Object Point Cloud Classification
We first evaluate the performance of our model on object point cloud classification tasks using the object expert encoder of Point-MoDE. We conduct point cloud classification on two of the most commonly used object point cloud datasets: ScanObjectNN Uy et al. (2019) and ModelNet40 Wu et al. (2015). ScanObjectNN contains 15K real scanned point clouds, each with various backgrounds, occlusions, and noise, which effectively assesses the model’s robustness. ModelNet40 includes 12K synthetic point clouds belonging to 40 different categories, with each point cloud being complete and clean, providing a better representation of 3D object shapes.
Following previous work Dong et al. (2023); Liang et al. (2024), we use 2K points as input for ScanObjectNN, apply simple rotation for data augmentation, and report classification accuracy without using voting. For ModelNet40, we use 1K points as input, apply scale and translate data augmentation, and report classification accuracy with the standard voting mechanism.
As presented in Table 1, firstly, compared to previous single-domain baselines (Point-MAE, Point-BERT, and PointGPT), our Point-MoDE method achieves consistent and stable improvements, indicating that leveraging scene domain knowledge to assist object domain learning is an effective strategy. This approach not only preserves object representations but also enables the learning of more comprehensive knowledge. Secondly, our Point-MoDE shows a significant improvement on the real-world scanned point cloud dataset ScanObjectNN, while the improvement on the synthetic point cloud dataset ModelNet40 is smaller. This is because we use randomly sampled scene data during pretraining, which lacks explicit semantics. Such non-semantic data is more similar to real scanned point clouds, thereby aiding the model in learning more realistic point cloud representations.
4.2.2 Object Point Cloud Part Segmentation
We assess the performance of our Point-MoDE in part segmentation using the ShapeNetPart dataset Chang et al. (2015), comprising 16,881 samples across 16 categories. We utilize the same segmentation setting after the pre-trained encoder as in previous works Pang et al. (2022); Zhang et al. (2022) for fair comparison. The experimental results displayed in Table 2. Our method also achieves consistent improvements compared to single-domain approaches, demonstrating the superiority of leveraging object knowledge to assist in scene tasks. This improvement appears less pronounced compared to detection tasks, which can be attributed to the inherently challenging nature of fine-grained tasks like segmentation. Therefore, even relatively small improvements are sufficient to highlight the effectiveness of our approach.
Methods Reference $\mathrm{mIoU}_{c}$ $\mathrm{mIoU}_{I}$
Supervised Learning Only
PointNet++ Qi et al. (2017b) NeurIPS 2017 81.9 85.1
PointMLP Ma et al. (2022) ICLR 2022 84.6 86.1
Single-Modal Self-Supervised Learning
PointContrast Xie et al. (2020) ECCV 2020 - 85.1
IDPT Zha et al. (2023b) ICCV 2023 83.8 85.9
Point-BERT Yu et al. (2022) CVPR 2022 84.1 85.6
MaskPoint Liu et al. (2022) ECCV 2022 84.4 86.0
Point-MAE Pang et al. (2022) ECCV 2022 84.2 86.1
ACT Dong et al. (2023) ICLR 2023 84.7 86.1
Point-FEMAE Zha et al. (2024a) AAAI 2024 84.9 86.3
PointGPT-B Chen et al. (2024) NeurIPS 2023 84.5 86.5
Point-MoDE w/ Point-BERT Ours 84.6($\uparrow$ 0.5) 86.2($\uparrow$ 0.6)
Point-MoDE w/ Point-MAE Ours 84.5($\uparrow$ 0.3) 86.3($\uparrow$ 0.2)
Point-MoDE w/ PointGPT-B Ours 84.9($\uparrow$ 0.4) 86.5(--)
@@ 4.2.3 Scene Point Cloud Detection
We further fine-tune the pre-trained Point-MoDE on scene-level object detection tasks. At this stage, we primarily rely on the scene expert encoder to process scene-level inputs. Simultaneously, we randomly select 32 point blocks from the scene point cloud and use the 32 shared object encoders to handle these local point blocks. During the decoding phase, we integrate the results from the scene expert encoder with random queries from the scene, helping the scene expert to better understand scene details. We use ScanNetV2 Dai et al. (2017), to evaluate our model’s scene understanding capabilities.
Table 3 shows our detection results, our Point-MoDE model has shown significant improvements compared with other single domain models. In particular, there is an improvement of 6-9% in the $AP_{50}$ metric. This substantial improvement is mainly attributed to our Point-MoDE using multiple object encoders to assist the scene encoder in analyzing the overall scene. This approach enables the scene model to focus on more local details, thereby enhancing scene understanding.
Methods Reference $AP_{25}$ $AP_{50}$
Supervised Learning Only
VoteNet Qi et al. (2019) ICCV 2019 58.6 33.5
3DETR Misra et al. (2021) ICCV 2021 62.1 37.9
Self-Supervised Learning
PointContrast Xie et al. (2020) ECCV 2020 58.5 38.0
DepthContrast Zhang et al. (2021) ICCV 2021 61.3 -
Point-BERT Yu et al. (2022) CVPR 2022 61.0 38.3
PiMAE Chen et al. (2023) CVPR 2023 62.6 39.4
Point-MAE Pang et al. (2022) ECCV 2022 59.5 41.2
ACT Dong et al. (2023) ICLR 2023 63.8 42.1
MaskPoint Liu et al. (2022) ECCV 2022 64.2 42.1
PointGPT-B Pang et al. (2022) NeurIPS 2023 63.1 42.8
PointDif Zheng et al. (2024) CVPR 2024 - 43.7
Point-MoDE w/ Point-BERT Ours 65.4($\uparrow$ 4.4) 47.1($\uparrow$ 8.8)
Point-MoDE w/ Point-MAE Ours 66.2($\uparrow$ 6.7) 48.8($\uparrow$ 7.6)
Point-MoDE w/ PointGPT-B Ours 66.7($\uparrow$ 3.6) 48.9($\uparrow$ 6.1)
@@
4.3 Scene Semantic Segmentation
@@
We have conducted experiments on scene-level semantic segmentation tasks to assess the performance of Point-MoDE in classifying each point in a scene into semantic categories. We validated our model using the indoor S3DIS Armeni et al. (2016) dataset for semantic segmentation tasks. Specifically, we tested the model on Area 5 while training on other areas and report the mean IoU (mIoU) and mean Accuracy (mAcc). To ensure a fair comparison, we used the same codebase based on the PointNext Qian et al. (2022a) baseline and employed identical decoders and semantic segmentation heads.
@@
The experimental results are shown in the table 4. Compared to training the PointNeXt model from scratch, our method improves the mIoU score by 2.3%. It also shows significant improvements over other pretraining models, such as Point-MAE Pang et al. (2022) and PointDif Zheng et al. (2024). This enhancement is largely due to our block-to-scene pretraining, which equips the model with strong scene understanding capabilities and further demonstrates the generalizability of our approach.
Methods Reference mIoU mAcc
Supervised Learning Only
PointNeXt Qian et al. (2022a) NeurIPS 2022 68.5 75.1
Pix4Point Qian et al. (2022b) 3DV 2024 69.6 75.2
Self-Supervised Learning
Point-BERT Yu et al. (2022) CVPR 2022 68.9 76.1
MaskPoint Liu et al. (2022) ECCV 2022 68.6 74.2
Point-MAE Pang et al. (2022) ECCV 2022 68.4 76.2
PointDif Zheng et al. (2024) CVPR 2024 70.0 77.1
PointGPT-B Pang et al. (2022) NeurIPS 2023 70.4 77.9
Point-MoDE w/ Point-BERT Ours 69.3($\uparrow$ 0.4) 76.8($\uparrow$ 0.7)
Point-MoDE w/ Point-MAE Ours 69.2($\uparrow$ 0.8) 76.6($\uparrow$ 0.4)
Point-MoDE w/ PointGPT-B Ours 70.9($\uparrow$ 0.5) 78.1($\uparrow$ 0.2)